# Hybrid physics-data-driven modeling for sea ice thermodynamics and transfer learning

### v1

@@ 1 Introduction
Over the past 15 years, machine learning (ML) has experienced exceptional growth, driven by remarkable progress in computational power and the pivotal introduction of GPU-accelerated training for deep neural networks as demonstrated with the breakthrough in image classification by krizhevsky2012imagenet with AlexNet. Since then, ML applications have permeated a wide range of scientific and technological domains, among which numerical weather and climate modeling has emerged as a prominent example, enabled by the extension of computer vision techniques to the geosciences. Recent advancements have enabled the development of global, fully data-driven, ML-based weather prediction (MLWP) systems (pathak2022fourcastnet; lam2023learning-graphcast; bi2023accurate-pangu; chen2025operational-fengwu; bodnar2025foundation). Trained primarily on ERA5 reanalysis data (hersbach2020era5), these models have shown comparable and sometimes superior skills with respect to ECMWF’s deterministic operational forecasting system, IFS-HRES, on short-to-medium range forecasts (rasp2024weatherbench). The main advantage of these models is their exceptional computational speed and efficiency, compared to classical physics-based models. However, they could remain limited by their lack of physical consistency and realism (Bonavita2024), which are crucial for meteorology in complex terrain, extreme event, coupled phenomena, as well as for seasonal and climate predictions. Additionally, the majority of them rely on high resolution physics-based models and data assimilation (DA) for their training and initialization. Another pragmatic approach, more conservative, is represented by hybrid modeling that combines ML with traditional physics-based models, in an attempt to extract the best of both while mitigating their individual weaknesses. This combination is often achieved through the replacement of physics-based or empirical parameterizations with data-driven ML components, with the aim of improving accuracy or reducing computational cost (chevallier1998neural; ogorman2018; rasp2018deep; Bolton2019). In the context of sea ice modeling, driscoll2024parameter trained Neural Networks (NNs) to emulate an advanced physics-based melt pond parameterization within the ice column model Icepack (elizabeth_hunke_2025_16422921), while Horvat2022 focused on wave-induced ice floe fracturing through a neural-network emulation of a computationally expensive super-parameterization.Despite continuous progress, numerical weather and climate model predictions inevitably remain affected by systematic errors, due to discretization errors, imperfect subgrid-scale parameterizations or inaccurate boundary conditions, initial conditions and forcings. Therefore, instead of focusing on specific process parameterizations, another research branch on hybrid modeling targeted the model error, developing ML-based model error parameterizations that complement physically derived models during execution (Watson2019). Similar to MLWP models, these hybrid models for model error correction leverage DA that, along with the increasing volume and quality of observational data, plays a fundamental role in geoscience, by providing optimal estimates (analyses) of the state of the earth system (Carrassi2018). In particular, analysis increments (differences between analysis and forecast states) are exploited to infer state-dependent systematic model errors (carrassi2011treatment). Key contributions along these lines include the work of brajard2021combining and Farchi2021, who trained NNs as additive integrated correction term between two forecast times, i.e. in the resolvent of geophysical models, learning from analysis increments, with particular focus on the role of observation noise and sparseness. Farchi2021comparison further investigated the potential of tendency correction and introduced an online learning strategy, in which both the model states and the model parameter are estimated at the same time as soon as new observations become available. Similar approaches to hybrid modeling have also been applied to more realistic scenarios, such as the work of Watt-Meyer2021, who trained a ML model to emulate nudging tendency terms within a weather forecast system; farchi2025 who learned a neural network correction to the Integrated Forecasting System of the ECMWF both offline and online; du2025reducing, who developed ML-based corrections for ocean mixed-layer temperature biases in a global climate model; and chapman2025improving, who improved multidecadal atmospheric predictions through online bias correction trained on nudging increments. This type of hybrid modeling has also been explored for sea ice. For example, Finn2023 employed convolutional U-nets to parametrize model error due to unresolved subgrid scale of sea-ice dynamics within an unstructured finite element model with a Maxwell elasto-brittle rheology. More recently, Gregory2024 successfully trained Convolutional Neural Networks (CNNs) to correct systematic errors of a global coupled ocean-sea ice model, demonstrating improvements achieved by iterating DA and ML steps (bocquet2020_fds). Building on this work, gregory2026advancing developed a hybrid global coupled ice–atmosphere–ocean climate model, highlighting the importance of fully coupled training for generalization in free forecasts. He2025 applied ML for online error correction within an earth system model for seasonal arctic sea ice prediction and compared it with an offline approach where ML post-processes and calibrates model predictions after the simulation.
@@ 3 Hybrid framework
The hybrid model developed in this work combines the physics-based Icepack model with an additive ML component that acts as a correction to the model resolvent. Trained on DA analysis increments, the ML component estimates systematic forecast biases at a prescribed lead time and iteratively corrects the model state. The rationale for this approach is that when the analyses are accurate enough to serve as proxies for the unknown true states (e.g., negligible analysis error), the corresponding analysis increments largely reflect forecast errors arising from model deficiencies (carrassi2011treatment). Under these assumptions, analysis increments provide suitable training targets for ML models that estimate state-dependent model error conditioned on forecast states (bocquet2020_fds; brajard2021combining; Farchi2021comparison; bocquet2023surrogate; Finn2023). The resulting hybrid approach is, therefore, reminiscent of a sequential DA cycle, with the ML corrections in place of analysis increments, as detailed in Section LABEL:ssec:hybrid_model.
In this study, training data are generated in an idealized setting with perfect, complete observations: under this assumption, the effect of initial condition error is eliminated. This intentional simplification allows us to focus on ML design and performance and to study in detail the strength and limitation of the proposed approach without the confounding effect derived from the presence of initial condition error and its interplay with the model error. The latter issue is addressed in the online hybrid model experiments (Section LABEL:sec:online_testing).
@@ 3.1 Training data generation
Given the assumption of perfect complete observations, for each location and model configuration, incorrect forecast trajectories $\bm{x}$ are produced by initializing models from ”true” reference full states $\bm{x}^{\textrm{t}}$ at $S^{\textrm{tr}}\!=522$ dates, $\{t_{s}\}$, spanning the period from 1 January 1994 to 31 December 2003, with a weekly stride $\tau_{w}=7$ days:
@@
where $\mathcal{M}_{t_{s}:t_{s}+\tau}$ denotes the resolvent of the physics-based model from time $t_{s}$ to $t_{s}+\tau$, and $\bm{F}_{t_{s}:t_{s}+\tau}$ the corresponding atmospheric forcings over the forecast window. Each simulation is run up to a maximum lead time $\tau=180$ days, with outputs stored every 5 days. The analysis of forecast error statistics as a function of lead time and initialization date (Section LABEL:sec:lead-sdata_errors, Supporting Information) led to selecting a fixed training lead time of $\tau_{\textrm{tr}}=60$ days, which provides a suitable balance between forecast degradation and signal strength. In more realistic settings, the forecast is expected to degrade faster, leading to a shorter training lead time. In general, a suitable balance should exist for any reasonable forecast model. Machine learning datasets are then constructed by pairing features with targets at the selected lead time $\tau_{\textrm{tr}}$. The targets are the forecast errors, $\bm{e}_{s}=\bm{x}_{t_{s}+\tau_{\textrm{tr}}}-\bm{x}^{\textrm{t}}_{t_{s}+\tau_{\textrm{tr}}},$ while the feature vectors, $\bm{z}_{s}=\mathcal{G}\!\left(\bm{x}_{t_{s}+\tau_{\textrm{tr}}},\;\bm{F}_{t_{s}+\tau_{\textrm{tr}}}\right),$ comprises model state variables, atmospheric forcings and diagnostics at time $t_{s}+\tau_{\textrm{tr}}$, where $\mathcal{G}$ selects and/or derives the predictors used as ML inputs. The resulting supervised learning dataset, combining data from all locations, for a single perturbed configuration is $\mathcal{D}=\big\{\left(\bm{z}_{i},\bm{e}_{i}\right)\big\}_{i=1}^{N},$ and counts $N=L\times S^{\textrm{tr}}=3132$ instances.
@@ 3.2 Machine learning regression task
The objective of the ML model is to approximate the mapping between input features $\bm{z}$ and the forecast error $\bm{e}$ through a NN, $\bm{g}(\bm{\theta},\bm{z})$, where $\bm{\theta}$ denotes the network trainable parameters. In this study, we focus on predicting the component-wise forecast errors in ice concentration, $\bm{a}=(a_{1},\dots,a_{C})$, and ice volume, $\bm{v}=(v_{1},\dots,v_{C})$. Two independent NNs, for ice concentration and ice volume, are implemented to prevent cross-task interference and ensure stable learning. Each network is trained by minimizing the following cost function,
$\mathcal{L}(\bm{\theta}^{\varphi})=\overline{\left(\bm{g}(\bm{\theta}^{\varphi},\bm{z}_{i})-\bm{e}^{\varphi}_{i}\right)^{2}},\qquad\varphi\in\{\bm{a},\bm{v}\}$ (2)
where $\overline{(\cdot)}$ denotes the average over the dataset instances and the five ice categories, and $\bm{e}^{\varphi}_{i}$ represents the forecast error, with $\varphi$ indicating either ice concentration or ice volume. The NN functional form $\bm{g}$ (i.e., its architecture) is identical for both ice concentration and volume; the distinction arises solely from the respective trained weights. The inputs and targets are normalized by per-variable means and standard deviations, estimated from the training dataset. Table LABEL:tab:feat_targ summarizes the target and features variables used. At inference time, the combined NN error prediction is given by
$\bm{g}\!\left(\hat{\bm{\theta}},\,\bm{z}\right)=\left[\bm{g}\!\left(\hat{\bm{\theta}}^{\bm{a}},\,\bm{z}\right),\bm{g}\!\left(\hat{\bm{\theta}}^{\bm{v}},\,\bm{z}\right)\right]$ (3)
where $\bm{g}\!\left(\hat{\bm{\theta}}^{\bm{a}},\,\bm{z}\right)$ and $\bm{g}\!\left(\hat{\bm{\theta}}^{\bm{v}},\,\bm{z}\right)$ denote the trained networks’ predictions for ice concentration and ice volume, respectively.

### v2

@@ 1 Introduction
Over the past 15 years, machine learning (ML) has experienced exceptional growth, driven by remarkable progress in computational power and the pivotal introduction of GPU-accelerated training for deep neural networks as demonstrated with the breakthrough in image classification by krizhevsky2012imagenet with AlexNet. Since then, ML applications have permeated a wide range of scientific and technological domains, among which numerical weather and climate modeling has emerged as a prominent example, enabled by the extension of computer vision techniques to the geosciences. Recent advancements have enabled the development of global, fully data-driven, ML-based weather prediction (MLWP) systems (pathak2022fourcastnet; lam2023learning-graphcast; bi2023accurate-pangu; chen2025operational-fengwu; bodnar2025foundation). Trained primarily on ERA5 reanalysis data (hersbach2020era5), these models have shown comparable and sometimes superior skills with respect to ECMWF’s deterministic operational forecasting system, IFS-HRES, on short-to-medium range forecasts (rasp2024weatherbench). The main advantage of these models is their exceptional computational speed and efficiency, compared to classical physics-based models. However, they could remain limited by their lack of physical consistency and realism (Bonavita2024), which are crucial for meteorology in complex terrain, extreme event, coupled phenomena, as well as for seasonal and climate predictions. Additionally, the majority of them rely on high resolution physics-based models and data assimilation (DA) for their training and initialization. Another pragmatic approach, more conservative, is represented by hybrid modeling that combines ML with traditional physics-based models, in an attempt to extract the best of both while mitigating their individual weaknesses. This combination is often achieved through the replacement of physics-based or empirical parameterizations with data-driven ML components, with the aim of improving accuracy or reducing computational cost (chevallier1998neural; ogorman2018; rasp2018deep; Bolton2019). More recently, kochkov2024neural introduced NeuralGCM, an end-to-end differentiable hybrid general atmospheric circulation model that combines a dynamical core with learned physics and achieves competitive performance for both weather forecasting and climate simulation. In the context of sea ice modeling, driscoll2024parameter trained Neural Networks (NNs) to emulate an advanced physics-based melt pond parameterization within the ice column model Icepack (elizabeth_hunke_2025_16422921), while Horvat2022 focused on wave-induced ice floe fracturing through a neural-network emulation of a computationally expensive super-parameterization.Despite continuous progress, numerical weather and climate model predictions inevitably remain affected by systematic errors, due to discretization errors, imperfect subgrid-scale parameterizations or inaccurate boundary conditions, initial conditions and forcings. Therefore, instead of focusing on specific process parameterizations, another research branch on hybrid modeling targeted the model error, developing ML-based model error parameterizations that complement physically derived models during execution (Watson2019). Similar to MLWP models, these hybrid models for model error correction leverage DA that, along with the increasing volume and quality of observational data, plays a fundamental role in geoscience, by providing optimal estimates (analyses) of the state of the earth system (Carrassi2018). In particular, analysis increments (differences between analysis and forecast states) are exploited to infer state-dependent systematic model errors (carrassi2011treatment). Key contributions along these lines include the work of brajard2021combining and Farchi2021, who trained NNs as additive integrated correction term between two forecast times, i.e. in the resolvent of geophysical models, learning from analysis increments, with particular focus on the role of observation noise and sparseness. Farchi2021comparison further investigated the potential of tendency correction and introduced an online learning strategy, in which both the model states and the model parameter are estimated at the same time as soon as new observations become available. Similar approaches to hybrid modeling have also been applied to more realistic scenarios, such as the work of Watt-Meyer2021, who trained a ML model to emulate nudging tendency terms within a weather forecast system; farchi2025 who learned a neural network correction to the Integrated Forecasting System of the ECMWF both offline and online; du2025reducing, who developed ML-based corrections for ocean mixed-layer temperature biases in a global climate model; and chapman2025improving, who improved multidecadal atmospheric predictions through online bias correction trained on nudging increments. This type of hybrid modeling has also been explored for sea ice. For example, Finn2023 employed convolutional U-nets to parametrize model error due to unresolved subgrid scale of sea-ice dynamics within an unstructured finite element model with a Maxwell elasto-brittle rheology. More recently, Gregory2024 successfully trained Convolutional Neural Networks (CNNs) to correct systematic errors of a global coupled ocean-sea ice model, demonstrating improvements achieved by iterating DA and ML steps (bocquet2020_fds). Building on this work, gregory2026advancing developed a hybrid global coupled ice–atmosphere–ocean climate model, highlighting the importance of fully coupled training for generalization in free forecasts. He2025 applied ML for online error correction within an earth system model for seasonal arctic sea ice prediction and compared it with an offline approach where ML post-processes and calibrates model predictions after the simulation.
@@ 3 Hybrid framework
The hybrid model developed in this work combines the physics-based Icepack model with an additive ML component that acts as a correction to the model resolvent. In the framework considered here, the ML component is trained on forecast errors computed with respect to reference true states at a prescribed lead time, and is then used to sequentially correct the model state. This corresponds to an idealized case of a more general approach in which DA analysis increments are used as training targets. The rationale for such an approach is that, when the analyses are accurate enough to serve as proxies for the unknown true states (e.g., negligible analysis error), the corresponding analysis increments largely reflect forecast errors arising from model deficiencies (carrassi2011treatment). Under these assumptions, analysis increments provide suitable training targets for ML models that estimate state-dependent model error conditioned on forecast states (bocquet2020_fds; brajard2021combining; Farchi2021; Farchi2021comparison; bocquet2023surrogate; Finn2023; farchi2025). The resulting hybrid approach is, therefore, reminiscent of a sequential DA cycle, with the ML corrections in place of analysis increments, as detailed in Section LABEL:ssec:hybrid_model. A related, but distinct, approach is proposed by peng2024hybrid who use deep learning within a DA framework for parameter-error estimation of a chaotic toy model, also in the presence of initial-condition errors.
In this study, training data are generated in an idealized setting with perfect, complete observations: under this assumption, DA is bypassed, and forecast errors relative to the reference true state are used directly as training targets, thereby eliminating the effect of initial-condition error. This intentional simplification allows us to focus on ML design and performance and to study in detail the strength and limitation of the proposed approach without the confounding effect derived from the presence of initial condition error and its interplay with the model error. The latter issue is addressed in the online hybrid model experiments (Section LABEL:sec:online_testing).
@@ 3.1 Training data generation
Given the assumption of perfect and complete observations, for each location and model configuration, incorrect forecast trajectories $\bm{x}$ are produced by initializing models from ”true” reference full states $\bm{x}^{\textrm{t}}$ at $S^{\textrm{tr}}\!=522$ dates, $\{t_{s}\}$, spanning the period from 1 January 1994 to 31 December 2003, with a weekly stride $\tau_{w}=7$ days:
@@
where $\mathcal{M}_{t_{s}:t_{s}+\tau}$ denotes the resolvent of the physics-based model from time $t_{s}$ to $t_{s}+\tau$, and $\bm{F}_{t_{s}:t_{s}+\tau}$ the corresponding atmospheric forcings over the forecast window.
Thus, $\bm{x}_{t_{s}+\tau}$ represents the forecast state at lead time $\tau$, initialized from the true state at time $t_{s}$. The corresponding forecast error $\bm{e}_{s,\tau}$ at lead time $\tau$ is defined as the difference between the forecast state and the true state at the same time: $\bm{e}_{s,\tau}=\bm{x}_{t_{s}+\tau}-\bm{x}^{\textrm{t}}_{t_{s}+\tau}.$
Each simulation is run up to a maximum lead time $\tau=180$ days, with outputs stored every 5 days. The analysis of forecast error statistics as a function of lead time and initialization date (Section LABEL:sec:lead-sdata_errors, Supporting Information) led to selecting a fixed training lead time of $\tau_{\textrm{tr}}=60$ days, which provides a suitable balance between forecast degradation and signal strength. In more realistic settings, the forecast is expected to degrade faster, leading to a shorter training lead time. In general, a suitable balance should exist for any reasonable forecast model. Machine learning datasets are then constructed by pairing features with targets at the selected lead time $\tau_{\textrm{tr}}$. The targets are the forecast errors, $\bm{e}_{s}=\bm{x}_{t_{s}+\tau_{\textrm{tr}}}-\bm{x}^{\textrm{t}}_{t_{s}+\tau_{\textrm{tr}}},$ while the feature vector, $\bm{z}_{s}=\mathcal{G}\!\left(\bm{x}_{t_{s}+\tau_{\textrm{tr}}},\;\bm{F}_{t_{s}+\tau_{\textrm{tr}}}\right),$ comprises model state variables, atmospheric forcings and diagnostics at time $t_{s}+\tau_{\textrm{tr}}$, where $\mathcal{G}$ selects and/or derives the predictors used as ML inputs. The resulting supervised learning dataset, combining data from all locations, for a single perturbed configuration is $\mathcal{D}=\big\{\left(\bm{z}_{i},\bm{e}_{i}\right)\big\}_{i=1}^{N},$ and counts $N=L\times S^{\textrm{tr}}=3132$ instances.
@@ 3.2 Machine learning regression task
The objective of the ML model is to approximate the mapping between input features $\bm{z}$ and the forecast error $\bm{e}$ through a NN, $g(\bm{\theta},\bm{z})$, where $\bm{\theta}$ denotes the network trainable parameters. In the following, we denote by $\hat{\bm{\theta}}$ the trained parameters obtained after optimization. In this study, we focus on predicting the component-wise forecast errors in ice concentration, $\bm{a}=(a_{1},\dots,a_{C})$, and ice volume, $\bm{v}=(v_{1},\dots,v_{C})$. Two independent NNs, for ice concentration and ice volume, are implemented to prevent cross-task interference and ensure stable learning. Each network is trained by minimizing the following cost function,
$\mathcal{L}(\bm{\theta}^{\varphi})=\overline{\left(g(\bm{\theta}^{\varphi},\bm{z}_{i})-\bm{e}^{\varphi}_{i}\right)^{2}},\qquad\varphi\in\{\bm{a},\bm{v}\}$ (2)
where $\overline{(\cdot)}$ denotes the average over the dataset instances and the five ice categories, and $\bm{e}^{\varphi}_{i}$ represents the forecast error, with $\varphi$ indicating either ice concentration or ice volume. The NN functional form $g$ (i.e., its architecture) is identical for both ice concentration and volume; the distinction arises solely from the respective trained weights $\bm{\theta}^{\bm{a}}$ and $\bm{\theta}^{\bm{v}}$. The inputs and targets are normalized by per-variable means and standard deviations, estimated from the training dataset. Table 3.2 summarizes the target and features variables used. At inference time, the combined NN error prediction is given by
$g\!\left(\hat{\bm{\theta}},\,\bm{z}\right)=\left[g\!\left(\hat{\bm{\theta}}^{\bm{a}},\,\bm{z}\right),g\!\left(\hat{\bm{\theta}}^{\bm{v}},\,\bm{z}\right)\right]$ (3)
where $g\!\left(\hat{\bm{\theta}}^{\bm{a}},\,\bm{z}\right)$ and $g\!\left(\hat{\bm{\theta}}^{\bm{v}},\,\bm{z}\right)$ denote the trained networks’ predictions for ice concentration and ice volume, respectively.
Table 2: Features and targets of NN model. Variables with dimension 5 are defined over the ITD categories used by default in Icepack.